OpenStudy (anonymous):

A cyclist riding at a velocity of 5 ms^-1 braked with uniform deceleration and travelled 3 m before she came to a stop. How long did she take to stop?

OpenStudy (anonymous):

The only problem is that we require acceleration. This can be found out by - v^2-u^2 = 2*a*s v=0 , u=5 m/s , s= 3 m This is actually work energy theorem where we have cancelled masses on both sides.